# Different physical and numerical sources of scatter in the $M_{\star}$-$M_{\mathrm{BH}}$ relation and their connection to galaxy evolution

### v1

@@ Abstract
Observations have established that the masses of supermassive black holes (SMBHs) correlate tightly with the stellar masses of their host galaxies, albeit with substantial scatter. The size of this scatter as a function of galaxy mass and redshift contains valuable information about the origin of SMBHs and the physical nature of their co-evolution with galaxies. In this work, we highlight this connection by studying the scatter in the $M_{\mathrm{BH}}-M_{\star}$ relation for massive galaxies in the Illustris, IllustrisTNG (TNG), and EAGLE cosmological simulations. We find that the scatter in TNG is significantly lower than in Illustris and EAGLE, reflecting their different BH feedback models. By performing various numerical experiments, we quantify different contributions to the scatter in the simulations, and also identify a suitably defined intrinsic scatter. The intrinsic scatter in Illustris and EAGLE is $\sim 0.3$ dex at $z=0$, and is dominated by variations from BH accretion, whereas the smaller scatter of TNG is rather dominated by hierarchical merging, suggesting that the massive galaxies in TNG are more tightly quenched. Variations in the BH seed mass can contribute to the scatter of the $M_{\rm BH}-M_{\star}$ relation as well, but whether this still plays a role at $z=0$ depends on the feedback model. Simulations with disabled AGN feedback produce much higher scatter for low-mass galaxies than seen in our cosmological simulations, demonstrating the crucial influence of feedback for determining the co-evolution of SMBHs and their host galaxies in this regime. In contrast, an important factor in reducing the scatter for massive galaxies is hierarchical merging of mostly quenched systems. Based on our results, we expect that the scatter in the $M_{\mathrm{BH}}-M_{\star}$ relation at high redshift could be particularly powerful in providing clues to the origin of SMBHs.
@@
pubyear: 2024pagerange: Different physical and numerical sources of scatter in the $M_{\star}$-$M_{\mathrm{BH}}$ relation and their connection to galaxy evolutionDifferent physical and numerical sources of scatter in the $M_{\star}$-$M_{\mathrm{BH}}$ relation and their connection to galaxy evolution
@@ 1 Introduction
There are two main explanations for the $M_{\mathrm{BH}}-M_{\star}$ relation. Some works suggest that the relation may simply arise as a regression to the mean when many mergers occur (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011; Anglés-Alcázar et al., 2013). Peng (2007) use a simple random-merging model to study the formation of the relation between the black hole (BH) mass $M_{\rm BH}$ and stellar mass $M_{\star}$. They find that the relation between $M_{\rm BH}$ and $M_{\star}$ will become tight, with a logarithmic slope of close to unity after experiencing several mergers, regardless of the initial distribution of the BH and stellar mass. Hirschmann et al. (2010) extend Peng’s work to different models and find that the conclusion still holds in the $\Lambda$CDM cosmogony. Jahnke & Macciò (2011) further study the origins of the scaling relation between $M_{\rm BH}$ and $M_{\star}$ within the $\Lambda$CDM model including also BH accretion and star formation, confirming that the scaling relation can be established by hierarchical merging, despite the presence of star formation and BH accretion.
Another, arguably more popular explanation for the tight correlation between $M_{\rm BH}$ and $M_{\star}$ is that there is a co-evolution of SMBHs and their host galaxies (Kormendy & Ho, 2013; Heckman & Best, 2014; Sun et al., 2015; Zhuang & Ho, 2023) in which they mutually influence each other in some way. Over the past decades, many theoretical works focused on self-regulated growth mechanisms of the stellar content of galaxies and their SMBHs in which feedback processes play a critical role (Di Matteo et al., 2005; Springel et al., 2005; Hopkins et al., 2006; Ciotti & Ostriker, 2007; Sijacki et al., 2007; Booth & Schaye, 2009; Debuhr et al., 2011; Gaspari et al., 2012; Li et al., 2015; Hopkins et al., 2016; Weinberger et al., 2018; Yuan et al., 2018; Zinger et al., 2020; Su et al., 2021; Weinberger et al., 2023). Early works like Di Matteo et al. (2005) proposed a physical picture where the $M_{\rm BH}-M_{\star}$ relation is a direct consequence of the self-regulation of BH accretion by AGN feedback, triggered by galaxy mergers. Other works like Silk & Rees (1998) and Zinger et al. (2020) propose a related scenario where the causal origin of the tight relation between SMBHs and their host galaxies originates is the competition between the cumulative energy released from SMBHs and the potential binding energy of their host halo.
Advances in cosmological hydrodynamical simulations have allowed studies of the relation between SMBHs and their host galaxies in the cosmological context, including a modeling of astrophysical processes like BH accretion and star formation. In current cosmological simulations, the relation between BHs and their galaxies has in fact become a baseline for testing the validity of the underlying galaxy formation model (Vogelsberger et al., 2014; Schaye et al., 2015; Weinberger et al., 2018; Davé et al., 2019; Pakmor et al., 2023, etc.). Habouzit et al. (2021) have studied the $M_{\rm BH}-M_{\star}$ relation in different recent cosmological simulations and found that all of them can roughly reproduce the relation observed for massive galaxies. Also, Li et al. (2020) analysed the properties of BHs and their host galaxies in Illustris and IllustrisTNG, finding that the $M_{\rm BH}-M_{\star}$ relation has a slight offset in these two simulations compared to observations, and is tighter as them as well.
@@
In cosmological simulations, the scatter in the $M_{\rm BH}-M_{\star}$ relation may not purely come from physical origins. Because the gravitational and hydrodynamical processes governing the evolution of galaxies are highly non-linear, simulations exhibit aspects of dynamical chaos (Keller et al., 2019; Genel et al., 2019; Borrow et al., 2023b). Extremely small variations in any resolution element can quickly grow and turn into sizable differences. This can happen even for identical initial conditions, because variations in numerical round-off from mathematically allowed re-orderings of the operations in parallel codes are sufficient to introduce such small differences. In addition, since many current cosmological simulations resort to stochastic models to address certain sub-grid physics, for example whether or not a star particle should form, or whether stellar or BH feedback is triggered, is decided based on drawing random numbers. The exact sequence of random numbers used then becomes a source of numerical variations, and can thus contribute scatter to the $M_{\rm BH}-M_{\star}$ relation that is of purely numerical origin.
Several works have demonstrated these numerical variations explicitly. For example, Keller et al. (2019) performed a set of cosmological zoom-in simulations with different galaxy formation models but identical initial conditions, and found that the scatter in the stellar masses due to numerical variations and stochastic modeling grows with a Lyapunov rate, but then saturates due to stellar feedback. Genel et al. (2019) examined properties of galaxies in the cosmological simulations carried out with the TNG physics model and code, and also found similar trends. Borrow et al. (2023b) investigated variations due to numerics and stochastic modeling in cosmological simulations with the EAGLE model and confirmed that the random variability should be taken into account when interpreting results for individual galaxies.
@@
The paper is organized as follow. In Section 2, we briefly introduce the Illustris, IllustrisTNG, EAGLE, and MillenniumTNG simulation projects. The scatter of the $M_{\rm BH}-M_{\star}$ relation for massive galaxies seen in these simulations, and the impact of galaxy selection criteria and numerical resolution is also briefly discussed. In Section 3, we introduce possible origins for the scatter and analyze the data from numerical toy experiments to constrain the numerical contributions to the scatter in the TNG model. We also apply the derived formalism to other cosmological simulations and estimate their intrinsic scatter and compare it with observations. In Section 4, we discuss further aspects of galaxy evolution from the viewpoint of the scatter in the $M_{\rm BH}-M_{\star}$ relation, in particular how this relates to galaxy quenching and to overmassive BHs at high redshift. Finally, we summarize and discuss our results in Section 5.
@@ 2.1.1 Illustris
The Illustris project is a cosmological simulation with box size equal to $106.5\,{\rm cMpc}$ that was carried out with the TreePM magneto-hydrodynamical moving-mesh code AREPO (Springel, 2010). The mass resolutions of dark matter (DM) particles and gaseous cells are $6.3\times 10^{6}\,{\rm M_{\odot}}$ and $1.3\times 10^{6}\,{\rm M_{\odot}}$, respectively. The subgrid models for BH seeding, accretion and feedback, and the star formation prescription and stellar feedback applied in Illustris are described in Vogelsberger et al. (2014). We concisely summarize these subgrid treatments here. For more details, please refer to the original paper.
Black holes are modelled as sink particles at the center of galaxies. BH seeds with mass $m_{\rm seed}=1.42\times 10^{5}\,{\rm M_{\odot}}$ are placed in all FOF halos with mass higher than $7.1\times 10^{10}~{}{\rm M_{\odot}}$ and which do not contain a BH yet. The BH accretion rate is estimated with the Bondi–Hoyle–Lyttleton formula subject to an Eddington limit:
@@ 2.1.2 IllustrisTNG
The IllustrisTNG project (hereafter TNG; Marinacci et al., 2018; Naiman et al., 2018; Pillepich et al., 2018b; Springel et al., 2018; Nelson et al., 2018) is a suite of cosmological simulations of galaxy formation also performed with AREPO. The TNG project employs three different box sizes: $51.7\,{\rm cMpc}$, $110.7\,{\rm cMpc}$, and $302.6\,{\rm cMpc}$ on a side, and the corresponding simulations are referred to as TNG50, TNG100, and TNG300, respectively. The TNG50 box was carried out at four different resolutions, while the TNG100 and TNG300 simulations have been computed for three different resolutions each. The mass resolutions of the DM particles in the different resolutions of TNG100 are $7.5\times 10^{6}\,{\rm M_{\odot}}$, $6\times 10^{7}\,{\rm M_{\odot}}$, and $4.8\times 10^{8}\,{\rm M_{\odot}}$, while the mass resolutions of the gases cells in the different resolutions of TNG100 are $1.4\times 10^{6}\,{\rm M_{\odot}}$, $1.1\times 10^{7}\,{\rm M_{\odot}}$, and $8.9\times 10^{7}\,{\rm M_{\odot}}$. Throughout this work, we mainly utilize the data of the TNG100 simulation with the highest resolution, allowing us to best compare with the other simulations, which have similar mass resolution. However, we will also utilize the two TNG100 realizations with lower resolution to investigate the robustness and resolution dependence of our conclusions.
The subgrid models for BHs and stars in TNG100 are quite similar to the ones in Illustris, but differ in a small number of crucial aspects. We here briefly describe these differences. For a more detailed description please refer to Weinberger et al. (2018) and Pillepich et al. (2018a).
@@ 2.1.3 EAGLE
The EAGLE project (Schaye et al., 2015; Crain et al., 2015) is a cosmological simulations with a box size of $100\,{\rm cMpc}$ performed with the TreePM smoothed particle hydrodynamics (SPH) formulation ANARCHY, implemented in the GADGET-3 code. The mass resolutions of the DM particles and gas particles are $9.7\times 10^{6}\,{\rm M_{\odot}}$ and $1.8\times 10^{6}\,{\rm M_{\odot}}$, respectively. For a full descriptions of the subgrid models for BHs and stars in the EAGLE simulation, please refer to Schaye et al. (2015). We here only offer a brief summary.
The BH seeding algorithm in EAGLE is similar to Illustris and TNG. BH-seeds with a mass $M_{\rm seed}=1.48\times 10^{5}\,{\rm M_{\odot}}$ are introduced in DM halos with mass $M_{\rm DM}>1.48\times 10^{10}\,{\rm M_{\odot}}$. The BH accretion rate is calculated with a modified Bondi-Hoyle-Lyttleton formula following Rosas-Guevara et al. (2015), after taking the angular momentum of the accreting gas into account:
@@
where $\bar{c_{\rm s}}$ is the kernel-weighted sound speed, $V_{\Phi}$ is the circular speed of the gas around the BH, and $C_{\rm visc}$ is a free parameter representing the viscosity of the accreting gas. The AGN feedback is purely thermal and based on Booth & Schaye (2009). In this model, the AGN feedback energy is injected around the BH with a feedback efficiency of 0.015, but the feedback energy is accumulated before it is injected until the deposited feedback energy can increase the temperature of the surrounding gas by $\Delta T\gtrsim 10^{8.5}{\rm K}$. This is done to avoid that the thermal energy injection can be rendered inefficient by numerical overcooling problems that can arise when the low density of the heated gas can (at least initially) not be spatially resolved.
The star formation model in EAGLE is based on Schaye & Dalla Vecchia (2008). Star formation will take place when the hydrogen number density of the gas particles is higher than $0.1(Z/0.002)^{-0.64}\,{\rm cm^{-3}}$. The star formation rate is then calculated by a pressure-dependent formula derived from a Kennicutt–Schmidt law. A power-law equation of state, $p\propto\rho^{4/3}$, is prescribed when the gas particles become star-forming. Feedback from supernova is injected with a stochastic model in the form of thermal energy based on Dalla Vecchia & Schaye (2012). Similar to the treatment of AGN feedback and in order to prevent numerical overcooling problems, the feedback energy is accumulated and only deposited once the energy release can cause a temperature increase of $\Delta T\gtrsim 10^{7.5}{\rm K}$ in the surrounding gas.
@@ 2.1.4 MillenniumTNG
The recent MillenniumTNG project (hereafter MTNG; Delgado et al., 2023; Contreras et al., 2023; Hadzhiyska et al., 2023a, b; Pakmor et al., 2023; Hernández-Aguayo et al., 2023; Bose et al., 2023; Kannan et al., 2023; Ferlito et al., 2023; Barrera et al., 2023) is an extension of IllustrisTNG with a box size of $740\,{\rm cMpc}$ in its flagship hydrodynamical model (the project also contains very large dark matter simulations not considered here). This volume is around 15 times larger than that of TNG300, allowing a much better representation of rarer massive systems. The mass resolution of DM particles and gas particles in MTNG is $1.7\times 10^{8}\,{\rm M_{\odot}}$ and $3.1\times 10^{7}\,{\rm M_{\odot}}$, respectively, which is comparable to TNG300.
@@
Figure 1: Upper panels: The data samples in MTNG from $z=2$ to $z=0$ based on the selection criteria $\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$. Lower panels: The “main sequence” samples in MTNG after discarding outliers with a GMM clustering algorithm.
Figure 2: The selected data samples in Illustris, TNG, EAGLE and MTNG from $z=3$ to $z=0$, as labeled, with the selection criteria $\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$. The dashed lines show the results of our power-law fits to the relations.

### v2

# Different physical and numerical sources of scatter in the $_{\mathrm{BH}}-MM_{\star}$ relation and their connection to galaxy evolution

@@ Abstract
Observations have established that the masses of supermassive black holes (SMBHs) correlate tightly with the stellar masses of their host galaxies, albeit with substantial scatter. The magnitude of this scatter as a function of galaxy mass and redshift contains valuable information about the origin of SMBHs and the physical nature of their co-evolution with galaxies. In this work, we highlight this connection by studying the scatter in the $M_{\mathrm{BH}}-M_{\star}$ relation for massive galaxies in the Illustris, TNG100, and EAGLE cosmological simulations. We find that TNG100 shows significantly lower scatter than Illustris and EAGLE, reflecting different BH feedback models. Using numerical experiments, we separate different contributions to the scatter, including an intrinsic component. At $z=0$, Illustris and EAGLE show $\sim 0.3$ dex intrinsic scatter dominated by BH accretion, while the smaller scatter in TNG100 is dominated by hierarchical merging, implying more tightly quenched massive galaxies. BH seed mass variations can add scatter, though their impact at $z=0$ depends on the feedback model. Without AGN feedback the scatter is much larger for low-mass galaxies ($\gtrsim 0.5$ dex for $\log M_{\star}<10^{10.5},{\rm M_{\odot}}$ at z=0z=0–3), underscoring the crucial role of feedback in SMBH–galaxy co-evolution. In contrast, hierarchical merging of quenched systems is the main factor reducing scatter for massive galaxies. Based on our results, we expect that the scatter in the $M_{\mathrm{BH}}-M_{\star}$ relation at high redshift could be particularly powerful in providing clues to the origin of SMBHs.
@@
pubyear: 2024pagerange: Different physical and numerical sources of scatter in the $_{\mathrm{BH}}-MM_{\star}$ relation and their connection to galaxy evolutionDifferent physical and numerical sources of scatter in the $_{\mathrm{BH}}-MM_{\star}$ relation and their connection to galaxy evolution
@@ 1 Introduction
[…]
Previous analyses of observations have also shown that the establishment of the $M_{\rm BH}$ and $M_{\star}$ relation cannot be accounted for by mergers alone (Lahav et al., 2011; Ginat et al., 2016). An arguably more popular explanation for the tight correlation between $M_{\rm BH}$ and $M_{\star}$ is that there is a co-evolution of SMBHs and their host galaxies (Kormendy & Ho, 2013; Heckman & Best, 2014; Sun et al., 2015; Zhuang & Ho, 2023) in which they mutually influence each other in some way. Over the past decades, many theoretical works focused on self-regulated growth mechanisms of the stellar content of galaxies and their SMBHs in which feedback processes play a critical role (Di Matteo et al., 2005; Hopkins et al., 2006; Booth & Schaye, 2009; Debuhr et al., 2011; Weinberger et al., 2018). Early works like Di Matteo et al. (2005) proposed a physical picture where the $M_{\rm BH}-M_{\star}$ relation is a direct consequence of the self-regulation of BH accretion by AGN feedback, triggered by galaxy mergers. Other works like Silk & Rees (1998) and Zinger et al. (2020) propose a related scenario where the causal origin of the tight relation between SMBHs and their host galaxies originates from the competition between the cumulative energy released from SMBHs and the potential binding energy of their host halo.
Advances in cosmological hydrodynamical simulations have allowed studies of the relation between SMBHs and their host galaxies in the cosmological context, including a modeling of astrophysical processes like BH accretion and star formation. In current cosmological simulations, the relation between BHs and their galaxies has in fact become a baseline for testing the validity of the underlying galaxy formation model (Vogelsberger et al., 2014; Schaye et al., 2015; Weinberger et al., 2018; Davé et al., 2019; Pakmor et al., 2023, etc.). Habouzit et al. (2021) have studied the $M_{\rm BH}-M_{\star}$ relation in different recent cosmological simulations and found that all of them can roughly reproduce the relation observed for massive galaxies. Also, Li et al. (2020) analysed the properties of BHs and their host galaxies in Illustris and IllustrisTNG, finding that the $M_{\rm BH}-M_{\star}$ relation has a slight offset in these two simulations compared to observations, and is also tighter than observed.
@@
In cosmological simulations, the scatter in the $M_{\rm BH}-M_{\star}$ relation may not purely come from physical origins. Because the gravitational and hydrodynamical processes governing the evolution of galaxies are highly non-linear, simulations exhibit aspects of dynamical chaos (Keller et al., 2019; Genel et al., 2019; Borrow et al., 2023b). Extremely small variations in any resolution element can quickly grow and turn into sizable differences. This can happen even for identical initial conditions, because variations in numerical round-off from mathematically allowed re-orderings of the operations in parallel codes are sufficient to introduce such small differences. In addition, many current cosmological simulations resort to stochastic models to address certain subgrid physics. For example, decisions such as whether a star particle should form, or whether stellar or BH feedback is triggered, are made based on drawing random numbers. The exact sequence of random numbers used then becomes a source of numerical variations, and can thus contribute scatter to the $M_{\rm BH}-M_{\star}$ relation that is of purely numerical origin.
Several works have demonstrated these numerical variations explicitly. For example, Keller et al. (2019) performed a set of cosmological zoom-in simulations with different galaxy formation models but identical initial conditions, and found that the scatter in the stellar masses due to numerical variations and stochastic modeling grows with a Lyapunov rate, but then saturates due to stellar feedback. Genel et al. (2019) examined properties of galaxies in the cosmological simulations carried out with the physics model and code of IllustrisTNG, and also found similar trends. Borrow et al. (2023b) investigated variations due to numerics and stochastic modeling in cosmological simulations with the EAGLE model and confirmed that the random variability should be taken into account when interpreting results for individual galaxies.
@@
The paper is organized as follow. In Section 2, we briefly introduce the Illustris, IllustrisTNG, EAGLE, and MillenniumTNG simulation projects. The scatter of the $M_{\rm BH}-M_{\star}$ relation for massive galaxies seen in these simulations, and the impact of galaxy selection criteria and numerical resolution is also briefly discussed. In Section 3, we introduce possible origins for the scatter and analyze the data from numerical toy experiments to constrain the numerical contributions to the scatter in the model of IllustrisTNG. We also apply the derived formalism to other cosmological simulations and estimate their intrinsic scatter and compare it with observations. In Section 4, we discuss further aspects of galaxy evolution from the viewpoint of the scatter in the $M_{\rm BH}-M_{\star}$ relation, in particular how this relates to galaxy quenching and to overmassive BHs at high redshift. Finally, we summarize and discuss our results in Section 5.
@@ 2.1.1 Illustris
The Illustris project is a cosmological simulation with box size equal to $106.5\,{\rm cMpc}$ that was carried out with the TreePM magneto-hydrodynamical moving-mesh code AREPO (Springel, 2010). The mass resolutions of dark matter (DM) particles and gaseous cells are $6.3\times 10^{6}\,{\rm M_{\odot}}$ and $1.3\times 10^{6}\,{\rm M_{\odot}}$, respectively, where the gas resolution corresponds to the target cell mass. The subgrid models for BH seeding, accretion and feedback, and the star formation prescription and stellar feedback applied in Illustris are described in Vogelsberger et al. (2014). We concisely summarize these subgrid treatments here. For more details, please refer to the original paper.
Black holes are modelled as sink particles at the center of galaxies. BH seeds with mass $m_{\rm seed}=1.42\times 10^{5}\,{\rm M_{\odot}}$ are placed in all FOF halos with mass higher than $7.1\times 10^{10}~{\rm M_{\odot}}$ and which do not contain a BH yet, by converting the densest gas cell in the halo into a BH particle. The BH accretion rate is estimated with the Bondi–Hoyle–Lyttleton formula subject to an Eddington limit:
@@ 2.1.2 IllustrisTNG
The IllustrisTNG project (hereafter TNG; Marinacci et al., 2018; Naiman et al., 2018; Pillepich et al., 2018b; Springel et al., 2018; Nelson et al., 2018) is a suite of cosmological simulations of galaxy formation also performed with AREPO. The TNG project employs three different box sizes: $51.7\,{\rm cMpc}$, $110.7\,{\rm cMpc}$, and $302.6\,{\rm cMpc}$ on a side, and the corresponding simulations are referred to as TNG50, TNG100, and TNG300, respectively. The TNG50 box was carried out at four different resolutions, while the TNG100 and TNG300 simulations have been computed for three different resolutions each. The mass resolutions of the DM particles in the different resolutions of TNG100 are $7.5\times 10^{6}\,{\rm M_{\odot}}$, $6\times 10^{7}\,{\rm M_{\odot}}$, and $4.8\times 10^{8}\,{\rm M_{\odot}}$, while the mass resolutions of the gases cells in the different resolutions of TNG100 are $4\times 10^{6}\,{\rm M_{\odot}}, 1.1$$1\times 10^{7}\,{\rm M_{\odot}}$, and $8.9\times 10^{7}\,{\rm M_{\odot}}$. Throughout this work, we mainly utilize the data of the TNG100 simulation with the highest resolution, allowing us to best compare with the other simulations, which have similar mass resolution. However, we will also utilize the two TNG100 realizations with lower resolution to investigate the robustness and resolution dependence of our conclusions.
The subgrid models for BHs and stars in TNG100 are quite similar to the ones in Illustris, but differ in a small number of crucial aspects. We briefly describe these differences here. For a more detailed description please refer to Weinberger et al. (2018) and Pillepich et al. (2018a).
@@ 2.1.3 EAGLE
The EAGLE project (Schaye et al., 2015; Crain et al., 2015) is a cosmological simulation with a box size of $100\,{\rm cMpc}$ performed with the TreePM smoothed particle hydrodynamics (SPH) formulation ANARCHY, implemented in the GADGET-3 code. The mass resolutions of the DM particles and gas particles are $9.7\times 10^{6}\,{\rm M_{\odot}}$ and $1.8\times 10^{6}\,{\rm M_{\odot}}$, respectively. For a full description of the subgrid models for BHs and stars in the EAGLE simulation, please refer to Schaye et al. (2015). We here only offer a brief summary.
The BH seeding algorithm in EAGLE is similar to Illustris and TNG. BH seeds with a mass $M_{\rm seed}=1.48\times 10^{5}\,{\rm M_{\odot}}$ are introduced in DM halos with mass $M_{\rm DM}>1.48\times 10^{10}\,{\rm M_{\odot}}$. The BH accretion rate is calculated with a modified Bondi-Hoyle-Lyttleton formula following Rosas-Guevara et al. (2015), after taking the angular momentum of the accreting gas into account:
@@
where $\bar{c_{\rm s}}$ is the kernel-weighted sound speed, $V_{\Phi}$ is the circular speed of the gas around the BH, and $C_{\rm visc}$ is a free parameter representing the viscosity of the accreting gas. The AGN feedback is purely thermal and based on Booth & Schaye (2009). In this model, the AGN feedback energy is injected around the BH with a feedback efficiency of 0.015, but the feedback energy is accumulated before it is injected until the deposited feedback energy can heat a single gas particle by at least $\Delta T\sim 10^{8.5}\,{\rm K}$. This is done to avoid that the thermal energy injection can be rendered inefficient by numerical overcooling problems that can arise when the low density of the heated gas can (at least initially) not be spatially resolved.
The star formation model in EAGLE is based on Schaye & Dalla Vecchia (2008). Star formation will take place when the hydrogen number density of the gas particles is higher than $0.1(Z/0.002)^{-0.64}\,{\rm cm^{-3}}$ and the gas temperature $T$ satisfies $\log T<\log T_{\rm EOS}+0.5$. The temperature floor, $T_{\rm EOS}$, follows a polytropic relation $P_{\rm EOS}\propto\rho^{4/3}$, normalized to $T_{\rm EOS}=8\times 10^{3}\,{\rm K}$ at $n_{\rm H}=0.1\,{\rm cm^{-3}}$. The star formation rate is then calculated by a pressure-dependent formula derived from a Kennicutt–Schmidt law. Feedback from core-collapse supernovae is implemented using a stochastic thermal injection scheme based on Dalla Vecchia & Schaye (2012). At an age of $3\times 10^{7}$ yr, each newly formed star particle probabilistically heats a subset of its neighbouring gas particles by a fixed temperature increase of $\Delta T\gtrsim 10^{7.5}\,{\rm K}$. This approach ensures that the energy injection is locally effective, avoiding artificial radiative losses due to numerical overcooling.
@@ 2.1.4 MillenniumTNG
The recent MillenniumTNG project (hereafter MTNG; see introductory and model descriptions in Pakmor et al., 2023; Hernández-Aguayo et al., 2023) is an extension of IllustrisTNG with a box size of $740\,{\rm cMpc}$ in its flagship hydrodynamical model (the project also contains very large dark matter simulations not considered here). Several papers have presented first scientific results based on MTNG, among them high-redshift galaxy populations (Kannan et al., 2023), large-scale galaxy clustering (Bose et al., 2023), weak lensing (Ferlito et al., 2023), and intrinsic alignments of galaxies (Delgado et al., 2023). The MTNG volume is around 300 times larger than that of TNG100, and 15 times larger than that of TNG300, allowing a much better representation of rarer massive systems. The mass resolution of DM particles and gas particles in MTNG is $1.7\times 10^{8}\,{\rm M_{\odot}}$ and $3.1\times 10^{7}\,{\rm M_{\odot}}$, respectively, which is comparable to TNG300, but worse than TNG100.
@@
Figure 1: Upper panels: The data samples in MTNG from $z=2$ to $z=0$ based on the selection criteria $\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$. Lower panels: The “main sequence” samples in MTNG after discarding outliers with a GMM clustering algorithm. The figure shows that GMM clustering can successfully remove the subpopulation that does not belong to the main relation. We will use the selected part shown in the lower panels for the analysis in the rest of this work.
Figure 2: The selected data samples in Illustris, TNG, EAGLE and MTNG from $z=3$ to $z=0$, as labeled, with the selection criteria $\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$. The dashed lines show the results of our power-law fits to the relations. The aim of this figure is to illustrate the distribution of selected data samples from the Illustris, TNG, EAGLE, and MTNG simulations, spanning redshifts from $z=3$ to $z=0$, with specific selection criteria ($\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$), and the corresponding fitting results.